OpenStudy (anonymous):

verify that cos (90 degrees +x)=-sin x

OpenStudy (ipwnbunnies):

I suggest to try some angles. 90 degrees, 30 degrees, 180 degrees, 60 degrees, etc. Do you know the trig functions of those special angles?

OpenStudy (anonymous):

i know some of them but i dont know which one to use haha

OpenStudy (anonymous):

wait wait i think i do

OpenStudy (anonymous):

cos (a-b)=cos a cos b+ sin a sin b

OpenStudy (ipwnbunnies):

No, like do you know cos(60) = sin(30) = 1/2?

OpenStudy (anonymous):

whot.

OpenStudy (ipwnbunnies):

Ugh, let's just use angles 90, 180, and 270 degrees. You know the trig functions of those, right?

OpenStudy (anonymous):

no ha im very confused deary

OpenStudy (ipwnbunnies):